OpenStudy (thefluffmuncher):

If f(x) varies directly with x, and f(x) = 56 when x = 8, find the value of f(x) when x = 2 A. 4 B. 7 C. 8 D. 14

OpenStudy (anonymous):

f(x) varies directly with x, i.e. f(x) a x --- Introducing an equal sign: "=" brings in a constant, k, i.e. f(x) = kx, f(x)/x = k When x = 8 & f(x) = 56, then k = 56/8 k = 7 => To find f(x), when x = 2 takes you back to the equation: f(x) = kx, where k = 7 & x = 2, then by substitution you'll have f(x) = 7 * 2 y = 14 (d).